and Applications Qs Gilberto buys only milk and cookies. In year 1, Gilberto earns $80, milk costs $2 per quart, and cookies cost $4 per dozen Use the green line (triangle symbol) to draw Gilbertos budget constraint on the following graph.
Answer #1

year / mix = 80-40 Cookies: 80 20 In year 2, New income- 80 +8= 88 milk priu- 2 + 0.2 = 2.2 lookies Piu - 4+0.4 = 4.4 New mil